OpenStudy (anonymous):

suppose that the nation demand and price for a certain type of energy-efficient exhaust fan are related by p=d(q)=462-5/7q, where p is the price in dollars and q is the demand in thousands. if the demand is for 189000 exhaust fans, than the price is?

OpenStudy (anonymous):

p = 462 - 5 q/ 7 or p = 462-5/7q ??

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

I'm asking you which one is correct from your post?